OpenStudy (anonymous):

pls help me solve this if domain of f is the set of all real numbers,the set of possible values of p is a. (-4,4) b. {4,-4} pls help me solve this if domain of f is the set of all real numbers,the set of possible values of p is a. (-4,4) b. {4,-4} @Mathematics

OpenStudy (aravindg):

?

OpenStudy (aravindg):

giv the function

OpenStudy (anonymous):

\[f(x)=\left(\begin{matrix}1 \\ \sqrt{x ^{2}-px+4}\end{matrix}\right)\]

OpenStudy (aravindg):

simple

OpenStudy (anonymous):

sry for the delay..

OpenStudy (aravindg):

u see its a cup shaped graph for a>0

OpenStudy (aravindg):

isnt 'it?

OpenStudy (anonymous):

not the graph...i cant find p's value